Define the multiplicative inverse in modular arithmetic and identify the multiplicative inverse of 6 mod 13 while explaining the algorithm used.



1
Expert's answer
2022-01-10T15:49:42-0500

The multiplicative inverse of a number x modulo m is a number y such that

"xy \equiv 1 \pmod m"

and we write "y \equiv x^{-1}"

The algorithm used for finding the inverse is known as Euclid's algorithm - it's a means of deducing the GCD of two given integers. In this case, we want to find x such that

"6x \equiv 1 \pmod{13}"

We have

"13 = 12 + 1 = 2\u20226 + 1"

(the remainder term here is the GCD)

Now,

"1 = 13- 2\u20226"

and taken mod 13, we have

"1 \equiv 13 - 2\cdot6 \equiv -2 \cdot 6 \pmod{13}"


"and \\

-2 \equiv 11 \pmod{13}"

so the inverse of 6 mod 13 is 11. Indeed,

"11\cdot6 \equiv 66 \equiv 65 + 1 \equiv 5\cdot13+1 \equiv 1\pmod{13}"